The driver of the train moving at 20 metre per second sites another train moving at 4 m per second on the same track and in the same direction he immediately applies break and history and begins to retire at 1 m per second square the minimum distance between the train for no collision is

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

u=20-4=16 m/s

a=1m/s/s

s=u*u/2a
s=16*16/2
= 128 m

Regards